# Hiding Local Manipulations on SAR Images: a Counter-Forensic Attack

### v1

@@ Abstract
The vast accessibility of Synthetic Aperture Radar (SAR) images through online portals has propelled the research across various fields. This widespread use and easy availability have unfortunately made SAR data susceptible to malicious alterations, such as local editing applied to the images for inserting or covering the presence of sensitive targets. Vulnerability is further emphasized by the fact that most SAR products, despite their original complex nature, are often released as amplitude-only information, allowing even inexperienced attackers to edit and easily alter the pixel content. To contrast malicious manipulations, in the last years the forensic community has begun to dig into the SAR manipulation issue, proposing detectors that effectively localize the tampering traces in amplitude images. Nonetheless, in this paper we demonstrate that an expert practitioner can exploit the complex nature of SAR data to obscure any signs of manipulation within a locally altered amplitude image. We refer to this approach as a counter-forensic attack. To achieve the concealment of manipulation traces, the attacker can simulate a re-acquisition of the manipulated scene by the SAR system that initially generated the pristine image. In doing so, the attacker can obscure any evidence of manipulation, making it appear as if the image was legitimately produced by the system. We assess the effectiveness of the proposed counter-forensic approach across diverse scenarios, examining various manipulation operations. The obtained results indicate that our devised attack successfully eliminates traces of manipulation, deceiving even the most advanced forensic detectors.
@@ I Introduction
Synthetic Aperture Radar (SAR) images are remote sensing data that use radar waves to create two-dimensional or three-dimensional representations of objects or landscapes. Over the last years, many online portals have given access to SAR images in easy-to-download and manageable data products. Among these portals, we can find the European Space Agency (ESA) Copernicus Open Access Hub [1], ICEYE [2] and Capella [3]. This phenomenon bolstered the research in the SAR field, with various applications in geology, hydrology, forestry, agriculture, urban planning, disaster management, intelligence, and military investigations [4, 5]. For example, SAR data can localize sensitive targets like army vehicles [6], ships [7], airports [8] and aircrafts [9].
Figure 1: High-level description of the paper goal. Given a manipulated SAR amplitude image (left), state-of-the-art SAR forensic detectors estimate a fingerprint (second column) that highligths local image inconsistencies. Our proposed counter-forensic attack simulates a SAR re-acquisition of the manipulated image. The attack leaves untouched the semantic content of the image (third column), but it conceals its tampering traces, such that the estimated fingerprint (right) does not allow anymore to localize the forgery. The manipulation area is sorrounded by the red countour.
@@
For this reason, SAR images are becoming objects of investigation by the multimedia forensic community [11, 10, 12]. Given the particular processing pipeline of SAR images, researchers have developed techniques tailored explicitly to them [11], proposing solutions for the localization of tampered-with pixel regions [10]. In particular, the authors of [10] have proposed a SAR forensic detector that uncovers traces of manipulations applied locally on SAR amplitude images. Given a SAR amplitude image, the detector processes it to estimate a fingerprint map that highlights local inconsistencies. If the map is homogeneous, the image is likely pristine; if the map shows a region with completely different characteristics than the rest, the image has been likely manipulated within that area. An example of the fingerprint returned by this detector is shown in Fig. 1. Given a manipulated image (left), the detector is able to localize the tampering area with high accuracy (second column).
@@
In this work, we prove that an experienced practitioner can exploit the complex nature of SAR data to conceal the manipulation traces from a locally manipulated SAR amplitude image. If any information on the SAR satellite used to acquire the original image is available, it can be used to simulate a complete re-acquisition of the manipulated scene. In doing so, we show that it is possible to produce an image with the same semantic content of the manipulated one, but with similar SAR characteristics to pristine ones, where the tampering traces have been hidden. A high-level example of the tackled goal is shown in Fig. 1. Given a manipulated amplitude SAR image (left), we produce a modified version of it (third column) such that the fingerprint estimated by the forensic detector does not highlight the forgery, but resembles that of a pristine image, hiding the manipulation artifacts (right).
The forensic community refers to this kind of operation as a counter-forensic attack. The goal of such attacks is usually twofold. On the one hand, the attack simulates the scenario in which a user produces a manipulation and is able to hinder the traces left by the tampering process. Notice that this does not involve malicious intents only; indeed, it can answer to privacy preserving needs in case specific image manipulations are inserted to hide sensitive targets. In this scenario, the attack allows to avoid the manipulation traces to be exposed. On the other hand, the attack allows to test the robustness of existing forensic detectors in spotting the manipulation even in presence of counter-forensic operations.
In our proposed counter-forensic attack, we alter the amplitude component of a complex SAR product, and we show that we are able to conceal the forensic footprints left by the manipulation by modeling the SAR acquisition and processing system with different strategies. In particular, we build a dataset of manipulated SAR amplitude images and validate the proposed pipeline in various scenarios, considering different manipulation operations and investigating in depth the effect of our methodology on their traces.
The achieved results show that the proposed attack wipes out the manipulation traces, fooling the most advanced forensic detectors. Moreover, the produced images carry typical SAR characteristics similar to original pristine ones, such that even an expert SAR analyst would not be able to realize the difference with the original data.
@@
  • We propose a counter-forensic attack that, leveraging the complex nature of SAR data, can fool advanced forensic detectors designed for SAR amplitude images.
  • The proposed attack enables to simulate scenarios in which a malicious user wants to delete their traces, but it allows as well to preserve the privacy in case sensitive targets must be hidden from the depicted scene.
  • The proposed attack enables to test the robustness and weaknesses of the existing forensic detectors in case of malicious attacks.
  • The manipulated images that undergo our proposed attack maintain similar characteristics than pristine data for what concerns their visual quality, SAR properties and frequency behaviour, thus rendering the attack difficult to be spot by expert researchers in the field.
@@ II-A2 Speckle pattern
As a consequence, the scene reflectivity image is composed by the contributions of manifold complex components, each due to a different scattering point on the target surface, showing a speckle pattern superimposed to the image of interest [14]. This pattern is maintained throughout the process of SAR data acquisition and focusing, thus it is reflected in the final reconstructed image $\mathbf{I}e^{j\mathbf{\Phi}}$. In Fig. 2 (right) we show a close-up of the previoulsy shown 2D SAR amplitude image, which is actually an SLC image thus it reasonaly contains a high amount of speckle noise. It is noticeable the effect of the speckle pattern as a granular noise superimposed to the scene of interest.
@@ II-B SAR amplitude image manipulations
Figure 3: Example of forgery creation process. From left to right: the donor image $\mathbf{I}_{D}$; the edited version of the donor image $E(\mathbf{I}_{D})$; the target image $\mathbf{I}_{T}$; the locally manipulated target image $\bar{\mathbf{I}}_{T}$. The splicing is performed by copying a region of the edited donor image into the target one. The manipulation area is sorrounded by the red countour.
@@ II-C SAR forensic detectors
Figure 4: Example of fingerprint $\bar{\mathbf{P}}$ and estimated tampering mask $\hat{\mathbf{M}}$ extracted by the detection method proposed in [10] from a spliced amplitude image $\bar{\mathbf{I}}$. The precise location of the spliced pixel area is reported in $\mathbf{M}$.
The forensic community has developed manifold image splicing detectors during the years, specifically focused on classic $8$-bit imagery like photographs or video frames [25, 26]. All the deployed detectors leverage the same assumption, that is, the acquisition device and processing operations leave peculiar traces on images, and these can be exploited to expose local forgeries  [27, 26].
The acquisition pipeline of SAR amplitude images is completely different than digital photographs or video sequences. Nonetheless, it has been shown that different SAR products reasonably contain different traces relative to the processing executed for generating them [10]. In this vein, the authors of [10] recently proposed a forensic detector for localizing splicing areas on SAR amplitude images. This detector highlights local inconsistencies between the donor and target pixel areas, and proves superior than state-of-the-art detectors on these specific images.
The detector training phase involves a siamese training of a denoising network architecture known as Denoising Convolutional Neural Network (DnCNN) [28]. The training process can be summarized by these three steps:
  1. 1.
    Consider a number of SAR amplitude images coming from different products.
  2. 2.
    Apply the DnCNN to patches extracted from different images to obtain a series of noise patterns.
  3. 3.
    Train the DnCNN to extract similar noise patterns for patches coming from the same product, and different patterns from patches coming from different products.
Eventually, the trained DnCNN is able to extract a noise-like fingerprint with the same size of the input image. This fingerprint contains traces related to the processing pipeline of the acquired product, highlighting potential splicing attacks as traces’ inconsistencies. Indeed, when analyzing pristine images, the fingerprint is self-consistent, whereas in case of spliced images it clearly highlights the edited regions.
More formally, let us consider a generic manipulated amplitude image as input to the detector. For simplicity, we omit the subscript $T$ from the manipulated images notation, taking for granted that all amplitude images indicated as $\bar{\mathbf{I}}$ have been manipulated in a specific target region of pixels, identified by the tampering mask $\mathbf{M}$. At first stage, the detector estimates a real-valued fingerprint map $\bar{\mathbf{P}}$, exposing local inconsistencies between pixels. Given $\bar{\mathbf{P}}$, the authors propose different methods to process $\bar{\mathbf{P}}$ for estimating a binary mask $\hat{\mathbf{M}}$ that identifies the forgery region. For an extensive description of the method, we refer the interested reader to [10].
Fig. 4 shows an example of fingerprint and tampering mask estimated by the detector. Notice that the splicing area is easily recognizable even from the fingerprint $\bar{\mathbf{P}}$. In general, the clearer the fingerprint artifacts, the better the estimation of the tampering mask.
@@ III Problem Formulation
Our goal is deleting the manipulation artifacts from $\bar{\mathbf{I}}$ such that forensic detectors cannot spot the manipulation traces anymore. In doing so, we do not want to deviate from neither the semantic content nor the visual quality of $\bar{\mathbf{I}}$. Moreover, we want to transfer to the attacked image the same properties of the pristine data acquired by the original SAR system, like the speckle-noise content and the traces left by the acquisition and processing operations. In other words, the final produced amplitude image must completely match the semantic characteristics of $\bar{\mathbf{I}}$ (i.e., it actually presents a forgered area), but it has to be transparent to visual inspection (i.e., the visual quality is mantained), to SAR properties inspection, and to the forensic detector, that must estimate it being pristine in its entire area. We define the attacked amplitude image as $\tilde{\mathbf{I}}$.
We consider as reference detector that proposed in [10], previously described in Section II-C, because it is specially tailored to SAR imagery and represents the state-of-the-art solution for spotting local image manipulations in the SAR field. Since this detector leverages the fingerprint extraction step as necessary operation for localizing the tampering area, our goal turns into concealing the forensic traces from $\tilde{\mathbf{I}}$ such that its fingerprint resembles that of of pristine image.
Fig. 5 reports the sketch of the tackled task: given a manipulated amplitude image $\bar{\mathbf{I}}$, the forensic detector returns the fingerprint $\bar{\mathbf{P}}$ that highlights some tampering traces in correspondence of the manipulated pixels. Our proposed method leverages information on the SAR system used to acquire the original data for performing a counter-forensic attack on $\bar{\mathbf{I}}$. The result is the amplitude image $\tilde{\mathbf{I}}$ which visually looks the same as $\bar{\mathbf{I}}$ and has the same SAR properties of a pristine image; however, if passed from the detector, the fingerprint $\tilde{\mathbf{P}}$ does not show any manipulation artifacts. We provide more details on the proposed counter-forensic attack in the next section.
@@ IV Proposed Methodology
Figure 6: Scheme of the proposed counter-forensic attack. By exploiting information on the original SAR system, we recreate the effect of a SAR acquisition through the same SAR system (or an estimate of it) used to acquire pristine data. To do so, we apply despeckling to the input manipulated image, then we simulate a re-acquisition by injecting speckle noise and filtering with the end-to-end SAR system. Finally, we match the histogram of the produced amplitude image with that of the manipulated amplitude image $\bar{\mathbf{I}}$.
@@
The effect of the proposed methodology is twofold. On the one hand, by acquiring the scene targets depicted in $\bar{\mathbf{I}}$ with the same acquisition system of the pristine data, we force the resulting amplitude image $\tilde{\mathbf{I}}$ to maintain similar SAR characteristics to the pristine data, like the amount of speckle noise and its spatial and frequency pattern. The final histogram matching completes the process by adjusting the pixel dynamics. On the other hand, the SAR acquisition processing chain acts as a counter-forensic attack that, by slightly modifying the pixel low-level information which are paramount for the detector, enables to hide the tampering traces. A detailed explanation of all the steps follows in the next lines.
@@ IV-B2 SAR System Simulation
    where $IF(\cdot)$ represents the 2D Inverse Fourier transform operator. The finally estimated 2D frequency response is defined as $\mathbf{H}_{D}=|F(\mathbf{h}_{D})|$. In this case, $\mathbf{H}_{D}$ satisfies the symmetry and positivity contraints.
@@ V Experimental Setup
In this section, we describe the setup of our experimental campaign. First, we present the employed SAR image dataset; second, we provide details on the forgery creation process; third, we give information on the counter-forensic attack parameters; finally, we describe the forensic detector setup.
@@ V-A Pristine Dataset
Figure 8: Examples of the considered SAR amplitude data. First column depicts pristine images. Second column shows manipulated version of the same images; the splicing area is contained inside the red box. Third column shows a close-up of the forgery area and of the sorrounding pixels (white box). Last column depicts a close-up with the proposed counter-forensic attack. Since the considered images have $16$-bit dynamics, we show their amplitudes in logarithmic scale.

### v2

@@ Abstract
The vast accessibility of Synthetic Aperture Radar (SAR) images through online portals has propelled the research across various fields. This widespread use and easy availability have unfortunately made SAR data susceptible to malicious alterations, such as local editing applied to the images for inserting or covering the presence of sensitive targets. To contrast malicious manipulations, in the last years the forensic community has begun to dig into the SAR manipulation issue, proposing detectors that effectively localize the tampering traces in amplitude images. Nonetheless, in this paper we demonstrate that an expert practitioner can exploit the complex nature of SAR data to obscure any signs of manipulation within a locally altered amplitude image. We refer to this approach as a counter-forensic attack. To achieve the concealment of manipulation traces, the attacker can simulate a re-acquisition of the manipulated scene by the SAR system that initially generated the pristine image. In doing so, the attacker can obscure any evidence of manipulation, making it appear as if the image was legitimately produced by the system. This attack has unique features that make it both highly generalizable and relatively easy to apply. First, it is a black-box attack, meaning it is not designed to deceive a specific forensic detector. Furthermore, it does not require a training phase and is not based on adversarial operations. We assess the effectiveness of the proposed counter-forensic approach across diverse scenarios, examining various manipulation operations. The obtained results indicate that our devised attack successfully eliminates traces of manipulation, deceiving even the most advanced forensic detectors.
@@ I Introduction
Synthetic Aperture Radar (SAR) images are remote sensing data that use radar waves to create two-dimensional or three-dimensional representations of objects or landscapes. Over the last years, many online portals have given access to SAR images in easy-to-download and manageable data products [1, 2, 3]. This phenomenon bolstered the research in the SAR field, with various applications in geology, hydrology, forestry, agriculture, urban planning, disaster management, intelligence, and military investigations [4, 5]. For example, SAR data can localize sensitive targets like army vehicles [6], ships [7], airports [8] and aircrafts [9].
Figure 1: High-level description of the paper goal. Given a manipulated SAR amplitude image (left), state-of-the-art SAR forensic detectors estimate a real-valued heatmap (second column) that highlights local image inconsistencies. Our proposed counter-forensic attack simulates a SAR re-acquisition of the manipulated image. The attack leaves untouched the semantic content of the image (third column), but it conceals its tampering traces, such that the estimated heatmap (right) does not allow anymore to localize the forgery. The manipulation area is surrounded by the red contour.
@@
For this reason, SAR images are becoming objects of investigation by the multimedia forensic community [11, 10, 12]. Given the particular processing pipeline of SAR images, researchers have developed techniques tailored explicitly to them [11], proposing solutions for the localization of tampered-with pixel regions [10]. An example of the localization heatmap returned by one state-of-the-art forensic detector is shown in Fig. 1. Given a manipulated image (left), the detector is able to localize the tampering area with high accuracy (second column).
@@
In this work, we prove that an experienced practitioner can exploit the complex nature of SAR data to conceal the manipulation traces from a locally manipulated SAR amplitude image. If any information on the SAR satellite used to acquire the original image is available, it can be used to simulate a complete re-acquisition of the manipulated scene. In doing so, we show that it is possible to produce an image with the same semantic content of the manipulated one, but with similar SAR characteristics to pristine ones, where the tampering traces have been hidden. A high-level example of the tackled goal is shown in Fig. 1. Given a manipulated amplitude SAR image (left), we produce a modified version of it (third column) such that the heatmap estimated by the forensic detector does not highlight the forgery, but resembles that of a pristine image, hiding the manipulation artifacts (right).
The forensic community refers to this kind of operation as a counter-forensic attack. The goal of such attacks is usually twofold. On the one hand, the attack simulates the scenario in which a user first performs a manipulation and then seeks to conceal the traces of it. Notice that this does not involve malicious intents only; indeed, it can answer to privacy preserving needs in case specific image manipulations are inserted to hide sensitive targets. In this scenario, the attack allows to avoid the manipulation traces to be exposed. On the other hand, the attack allows to test the robustness of existing forensic detectors in spotting the manipulation even in presence of counter-forensic operations.
In the literature, the counter-forensic attacks developed to fool tampering localization detectors are all tailored to standard digital pictures [15, 16, 17, 18, 19]. They often rely on the knowledge of the targeted detector’s functioning [15, 16, 17, 18] and can require access to the distribution of pristine samples [16, 18]. Moreover, when considering deep-learning-based methods, these techniques require training on datasets with characteristics similar to those of test data [15, 16]. Our proposed attack differs sensibly from these methodologies: first, it consists of a black-box attack, i.e., it is not tailored to deceive a specific forensic detector; second, it does not require a dataset of pristine and/or manipulated samples for training or performing adversarial operations.
In our proposed counter-forensic attack, we alter the amplitude component of a complex SAR product, and we show that we are able to conceal the forensic footprints left by the manipulation by modeling the SAR acquisition and processing system with different strategies. We purposely avoid any training pipeline, as we aim at simulating a scenario in which an attacker may have little knowledge of the SAR system used for acquiring the image they want to attack.
We build a dataset of manipulated SAR amplitude images and validate the proposed pipeline in various scenarios, considering different manipulation operations and investigating in depth the effect of our methodology on their traces. The achieved results show that the proposed attack wipes out the manipulation traces, fooling the most advanced forensic detectors and outperforming other counter-forensic solutions. Moreover, the produced images carry typical SAR characteristics similar to original pristine ones.
@@
  • We propose an attack that successfully targets advanced forensic detectors in a black-box scenario, meaning that it is not designed to deceive a specific detector, but it works generally to conceal the tampering traces.
  • The proposed attack does not require any training phase and does not depend on the access to a pristine dataset for calibration.
  • The attacker can rely on the single original complex image to achieve outstanding anti-forensic performance.
  • The manipulated images that undergo our proposed attack maintain similar characteristics than pristine data for what concerns their visual quality, SAR properties and frequency behaviour, thus rendering the attack difficult to be spotted.
@@
  • The proposed attack outperforms another counter-forensic solution lately proposed to conceal local manipulation traces [19].
@@ II-A2 Speckle pattern
As a consequence, the scene reflectivity image is composed of the contributions of manifold complex components, each due to a different scattering point on the target surface, showing a speckle pattern superimposed to the image of interest [14]. This pattern is maintained throughout the process of SAR data acquisition and focusing, thus it is reflected in the final reconstructed image $\mathbf{I}e^{j\mathbf{\Phi}}$. In Fig. 2 (right) we show a close-up of the previously shown 2D SAR amplitude image, which is actually an SLC image, thus it contains a reasonable amount of speckle noise. The effect of the speckle pattern is noticeable as a granular noise superimposed to the scene of interest.
@@ II-B SAR amplitude image manipulations
Figure 3: Example of forgery creation process. From left to right: the donor image $\mathbf{I}_{D}$; the edited version of the donor image $E(\mathbf{I}_{D})$; the target image $\mathbf{I}_{T}$; the locally manipulated target image $\bar{\mathbf{I}}_{T}$. The splicing is performed by copying a region of the edited donor image into the target one. The manipulation area is surrounded by the red contour.
@@ II-C SAR forensic detectors
The forensic community has developed manifold image splicing detectors during the years, specifically focused on classic $8$-bit imagery like photographs or video frames [30, 31, 32]. All the deployed detectors leverage the same assumption, that is, the acquisition device and processing operations leave peculiar traces on images, and these can be exploited to expose local forgeries [33, 31].
The acquisition pipeline of SAR amplitude images is completely different than digital photographs or video sequences. Nonetheless, it has been shown that different SAR products reasonably contain different traces relative to the processing executed for generating them [10]. In this vein, the authors of [10] recently proposed two forensic detectors, namely the SAR Adapted Extractor (SAE) and the Augmented SAR Adapted Extractor (ASAE), for localizing splicing areas on SAR amplitude images. Following a similar approach, the TruFor detector [32], which has shown outstanding performances on natural images and represents now the state of the art on them [34], can be adapted to SAR amplitude images, provided that the training pipeline is specifically designed to account for the unique characteristics of SAR data. We refer to this SAR-adapted version of the detector as TruForSAR (TruForS)  [^1] More details are reported in the supplemental materials..
These detectors highlight local inconsistencies between the donor and target pixel areas. More formally, let us consider a generic manipulated amplitude image as input to the detector. For simplicity, we omit the subscript $T$ from the manipulated images notation, taking for granted that all amplitude images indicated as $\bar{\mathbf{I}}$ have been manipulated in a specific target region of pixels, identified by the tampering mask $\mathbf{M}$. The three detectors estimate a real-valued heatmap $\bar{\mathbf{M}}$, ranging in $[0,1]$, that identifies the forgery region. Examples of the tampering heatmap $\bar{\mathbf{M}}$ estimated by the TruForS detector can be found in Figs. 1 and 4.
@@ III Problem Formulation
Our goal is deleting the manipulation artifacts from $\bar{\mathbf{I}}$ such that forensic detectors cannot spot the manipulation traces anymore. In doing so, we do not want to deviate from neither the semantic content nor the visual quality of $\bar{\mathbf{I}}$. Moreover, we want to transfer to the attacked image the same properties of the pristine data acquired by the original SAR system, like the speckle noise and the traces left by the acquisition and processing operations. In other words, the final produced amplitude image must completely match the semantic characteristics of $\bar{\mathbf{I}}$ (i.e., it actually presents a forged area), but it has to be transparent to visual inspection (i.e., the visual quality is maintained), to SAR properties inspection, and to the forensic detector, that must estimate it being pristine in its entire area. We define the attacked amplitude image as $\tilde{\mathbf{I}}$.
We consider as reference detectors those listed in Section II-C, because they are specially tailored to SAR imagery and represent the state-of-the-art solutions for spotting local image manipulations in the SAR field. Our goal turns into concealing the forensic traces from $\bar{\mathbf{I}}$ such that its estimated heatmap resembles that of a pristine image.
Fig. 4 reports the sketch of the tackled task: given a manipulated amplitude image $\bar{\mathbf{I}}$, the forensic detectors return a real valued heatmap $\bar{\mathbf{M}}$ that highlights some tampering traces in correspondence of the manipulated pixels. Our proposed method leverages information on the SAR system used to acquire the original data for performing a counter-forensic attack on $\bar{\mathbf{I}}$. The result is the amplitude image $\tilde{\mathbf{I}}$ which visually looks the same as $\bar{\mathbf{I}}$ and has the same SAR properties of a pristine image; however, if passed from the detectors, the heatmap $\tilde{\mathbf{M}}$ does not show any manipulation artifacts. We provide more details on the proposed counter-forensic attack in the next section.
@@ IV Proposed Methodology
Figure 5: Scheme of the proposed counter-forensic attack. By exploiting information of the original SAR system, we recreate the effect of a SAR acquisition through the same SAR system (or an estimate of it) used to acquire pristine data. To do so, we apply despeckling to the input manipulated image, then we simulate a re-acquisition by injecting speckle noise and filtering with the end-to-end SAR system. Finally, we match the histogram of the produced amplitude image with that of the manipulated amplitude image $\bar{\mathbf{I}}$.
@@
The effect of the proposed methodology is twofold. On the one hand, by acquiring the scene targets depicted in $\bar{\mathbf{I}}$ with the same acquisition system of the pristine data, we force the resulting amplitude image $\tilde{\mathbf{I}}$ to maintain similar SAR characteristics to the pristine data, like the amount of speckle noise and its spatial and frequency pattern. The final histogram matching completes the process by adjusting the pixel dynamics. On the other hand, the SAR acquisition processing chain acts as a counter-forensic attack that, by slightly modifying the pixel low-level information which are paramount for the forensic detectors, enables to hide the tampering traces. A detailed explanation of all the steps follows in the next lines.
@@ IV-B2 SAR System Simulation
    where $IF(\cdot)$ represents the 2D Inverse Fourier transform operator. The finally estimated 2D frequency response is defined as $\mathbf{H}_{D}=|F(\mathbf{h}_{D})|$. In this case, $\mathbf{H}_{D}$ satisfies the symmetry and positivity constraints.
@@ V Experimental Setup
In this section, we describe the setup of our experimental campaign. First, we present the employed SAR image dataset; second, we provide details on the forgery creation process; third, we give information on the counter-forensic attack parameters; finally, we describe the forensic detectors setup.